OpenStudy (anonymous):

6-9w=24 And a special thanks to Wombat:)

OpenStudy (anonymous):

6-9w=24 -9w=18 w=-2

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Thankyou wombat you are so sweet im awarding you another best answer:)

OpenStudy (anonymous):

Haha I'm glad I can help :)))

OpenStudy (anonymous):

Oh you did:D They should give you people award hahaha

OpenStudy (anonymous):

I like the sound of that XD

OpenStudy (anonymous):

Hahah yeah, i would vote for you:)

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

I feel bad about you answering all the questions because i dont want you to feel like im using you cuz ssure you im not, oh and i fanned you already:)

OpenStudy (anonymous):

Hhaha truly don't worry about it XD I love math and I am more than happy to help, especially since you have to get caught up. Its so hard to get caught up for me too. Miss a day for an event, BAM homework galore haha. SO to answer that...you're not using me :) I'm volunteering

OpenStudy (anonymous):

Haha yeah, your so lucky your smart at math, thank youso much for all your doing, yeah i hate getting caught up after a long missing week or something

OpenStudy (anonymous):

Well it balances out...im good with the "right or wrong answer" stuff like math and science, but i stink at the "subjective" and "more-right/more-wrong" stuff in english XD

OpenStudy (anonymous):

Haha really? I am really good in English and all the easy subjects but im terrible at math and science

OpenStudy (anonymous):

WOAAAH!! haha we are polar opposites :P

OpenStudy (anonymous):

Haha yeah